OpenStudy (anonymous):

Solve the equation 2^(3x) = 8^(4x/1/2) for x

OpenStudy (anonymous):

\[2^{3x}=8^{4x-\frac{ 1 }{ 2 }}\]

OpenStudy (asnaseer):

hint: use the fact that \(8=2^3\)

OpenStudy (anonymous):

I don't really see how that helps...

OpenStudy (asnaseer):

if you use that hint then you will effectively have something like:\[2^x=2^y\]which means you can state that:\[x=y\]

OpenStudy (asnaseer):

does that make sense now?

OpenStudy (lgbasallote):

that helps more than you'll ever know....it's the meaning of life...

OpenStudy (asnaseer):

:D

OpenStudy (anonymous):

3x = 12x - 3/2 9x = 3/2 x = 3/18 x = 1/6

OpenStudy (anonymous):

all done

OpenStudy (anonymous):

Wait! I wrote the equation wrong! It's 2^(3x) = 8*4^(x/1/2)

OpenStudy (asnaseer):

use the same hint

OpenStudy (asnaseer):

but in addition to that use: \(4=2^2\)

OpenStudy (anonymous):

Thanks!

OpenStudy (asnaseer):

yw and I'm glad you finally worked it out! :)